how many grams of ammonium nitrate need to be added to 500g of water so that the freezing point becomes -8.3 degree celcius
1
Expert's answer
2017-02-20T11:27:14-0500
Δtfreezing = K*m´
m´= Δtfreezing /K
m´= 8.3/1.86=4.46 Mol/kg
₰(NH₄NO₃ )= m´*m(H₂O )= 4.46Mol/kg*0.5kg = 2.23Mol
m (NH₄NO₃) = 2.23Mol*80g/Mol = 178.4g